Bent Gregorian Technical Observing with LUCIFER/AGw:
UT 2009 December 05

Observer: DThompson (@LBT), JHill (Tucson, first half)
Telescope Operator: ACeranski
SW support: PGrenz, NCushing

Summary

  • Gathered twilight flats data in J, K, Br_gam (unknown quality)
  • Investigated early-night collimation problems (refer IT#2444). Switched to DG coll.model, worked better.
  • Gathered short sample of field aberration measurements
  • Problems with oacserver failing (see IT#2446) lost 1:15 to this
  • Problems with primary failing to make chase - HP2 motor turning freely, HP not moving. (IT#2448) lost 3:15 to this

  • SDT Observations (clear, excellent seeing, telescope seems to be working)
    • Completed germeroth w/photom.std
    • Took imaging data for zibetti/PSB04, w/photom.stds
    • Attempted spectrum on PSB04, but MOS problem 0:52 lost

Preparations

Details

0:03 UT set up /data/luci/20091205/luci.20091205.* Bias frame taken (0001), saved both locally and to archive.

Twilight flats
00:13 Open, at zenith. Trying flats by hand (0002 = test in Br_gam + K).
0003 - 12 Br_gam + K (20 - 14k) 10s DCR
0013 - 22 Br_gam + clear (13 - 9k) 10s DCR
0023 - 32 K + clear (15k) 2s DCR
0033 - 42 Br_gam + clear (6 - 5k) 10s DCR
0043 - 52 K + clear (9 - 7k) 2s DCR
0053 - 55 J + clear () 2s DCR - saturated. Aborted reads and Inserted blind filter
0056 - 65 J + clear (6.5- 5k) 2s DCR
0066 - 75 J + clear (1k) 2s DCR after waiting a few minutes

Sunset: clear except for a few clouds around horizon. Outside=-1.5C M1=-7.1C RH=4%

00:46 Preset to ACT0185 for pointing correction. Br_gam in. Star w/lots of coma at 859, 776 (luci#0076). Problems were with Y collimation last night, currently we are at 1.6, offsetting to 0.6 (luci#0077 - looks better, and closer to rotcen). Offset another -1mm, better. Offset -0.5mm (now at -2.6 on global offsets for M1Y). Still better. Z offset +0.3...better. Another +0.3...better. Now at 885, 1248 (rotcen 1014, 1043). IE=-86. CA=-6. The GlobalOffsets put in for M1 were in the direction of those in use for direct Gregorian.

01:11 Preset to BS9186 for collimation. Converging.
0083 2s DCR on BS9186 in Br_gam. Send offset...M1 hit -Y limit. Add +2mm global offset in both M1 and M2, resend preset.

01:33 Experiment suggested by John: copy over the DG collimation model to LFBG and re-initialize everything. Replaced June 25 LFBG coll model with Oct 27 IRTC DG coll model. Send preset....initial collimation was better, but we hit a limit on offset. John talking to Andrew about the 5mm offset. New preset and the collimation looked OK with the caveat that the active optics correct translated into M2 Offload was X=-0.067 Y=-3.512 Z=0.218 RX=-38.4 RY=+1.1. (This is an order of magnitude larger than it is supposed to be. And it is similar to last night's result (IT#2444) even though the temperature gradient is reversed. Tonight we are warming the telescope and mirrors rather than cooling.) We build these corrections into the M1 Global Offsets which are now: X=-0.622 Y=-3.112 Z=1.050 RX=-88.3 RY=+5.9. The M2 Global Offsets were shifted in Y to recenter the primary: X=-7.010 Y=+2.121 Z=0 RX=+148.65 RY=+181.8.
Now the question is whether whatever caused this shift did it in a way to perturb the field aberration alignment of M2?

02:08 Check pointing model. Preset to ACT0204 (23h +0d) (at 1089, 988), IE=+8.6 CA=-15.
02:14 ACT0223 (0h +4d) at 1001, 1040
02:16 ACT0405 (0h +75d) at 1005, 1022
02:22 ACT0304 (20h +28d) at 1011, 1044
02:28 ACT0297 (3h +27d) at 995, 988
02:30 ACT0293 (0h +25d) at 1011, 1100
So, checked NSEW in the 45-60d EL range and one near zenith (+82d) and pointing looks good. The 5 positions after IE and CA corrected are within 7 arcsec of each other. Luci images through 0098.

02:33 Preset to BS9101 for field aberrations, but first: luci#0099 6x5s K integrated on BS9101. luci#0100 offset 30 +X for skysub. FWHM in K in 30s is ~4.4 pixels, 0.53 arcsec. No obvious field aberrations.

02:45 Turned over to John to gather a set of field aberration data.

02:46 cl < home$field_BS9101.cl (but, forgot to prepdir the /OldRepository/AGW_Data/091205/ directory)
02:48 cl < home$field_BS9101.cl and took wfsc 0002-0006, but IDL routines didn't process.
Realize that script is still set for LDG. Fix that by adding -f LFBG to the calls to my Python wrappers.
02:59 cl < home$field_BS9101.cl but then realize that IDL is trying to authorize as IRTC and not LUCIFER
Change authorization to IRTC, but then pointing is off.
03:00 IRAF error. Stop and restart IRAF window.
Copy the new LUCIFER collimation and pointing models into IRTCF so both are the same.
Adjust IE=1.0 CA=-16.3 as LUCIFER and then switch back to IRTC.
03:28 cl < home$field_BS9101.cl - But IDL still fails to send WF correction
Change IRAF script to use GCS ACTIVE correction on-axis.
03:33 cl < home$field_BS9101.cl - Seeing is now 0.55 arcsec.
03:37 preset to GS7 off-axis. But script gets syntax error.
Apparently it needed quotes around LFBG in the python call inside the IRAF script.
03:39 cl < home$field_BS9101.cl - uparm pipe error. Restart IRAF window again
03:41 cl < home$field_BS9101.cl
03:47 Gs7 again but IDL fails to process the off-axis image.
Called Doug. He fixed a bug that he had introduced this afternoon processing some old data.
Set Active Mode to Primary rather than Split.
IDL off-axis data is in /OldRepository/AGW_Data/091205/, and on-axis GCS WFS data is in /Repository/20091205/GCS/.
03:59 cl < home$field_BS9101.cl
04:00 on-axis
04:02 GS7 PA=15.41 Doug=-341 wfsc 0016-0018 Focus: -77, -83, +68 nm
on-axis
04:09 GS8 PA=124.74 Doug=-448 wfsc 0019-0021 Focus: -79, +95, +66
on-axis
04:16 GS4 PA=336.91 but wfsc 00022-24 are JUNK since guidestar seems to have been a galaxy
04:19 AGW3: Error on UMAC - couldn't move probe back to the on-axis position
04:20 stop GCSL, stopAGW -u 3 gives RPC error
04:22 stop and restart oacsrv
04:29 try to home probe - Call failed while checking error bits on UMAC.
Hand phone and telescope back to Aaron to call for software help.
05:19 Norm and friends have unscrambled the oacserver.
05:19 pointto BS9101 ACTIVE gs=4 (confirm it is very faint. maybe a galaxy?)
05:21 pointo BS9101 ACTIVE gs=3 - seeing is 0.5 arcsec at EL=45 deg
05:29 cl < home$field_BS9101b.cl (which only does GS3, and a check on GS=7)
05:30 on-axis
05:33 GS3 PA=307.27 Doug=-255 wfsc 0023-25 Focus: +149, +234, +238 nm
05:36 back on-axis
05:39 GS7 PA=15.41 Doug=-322 wfsc0026-28 Focus: -12, +51 +112 nm

### Seeing is good, executing germeroth SDT program 06:05 BS9105 J 5x2s luci#0120-0124
06:10 execute germeroth script on SDSS J02222-0003. Guider sees 0.8" at start. First set 5.1 pix FWHM = 0.61 arcsec. Files luci#0125-0157
06:53 BS9105 J 5x2s luci#0158-0159 TRACK mode - did not "set new value"

BS9122 (Zibetti PSB04 standard).
07:12 script Zibetti_PSB04_STD_ImgJ.txt failed to run.
07:13 BS9122 from my list. Probe is ~1 arcmin off center, vignettes.

Preset back to BS9122 using GS20 (coords inserted in UI by hand). Star is R2=15.79 mag. GCS acquisition image is 00035. Check flux etc in guider images in this sequence.

07:17 Getting Error while trying to compensate for flexure. Started 07:16:09 in LUCIFER log. Set flexure compensation off in UI and the error messages stopped.

07:27 BS9122 5x2s J PA=225. Guide probe still on field. Files luci#0167 - 0171

07:33 Execute Zibetti_PSB04_STD_ImgJ.txt. It was failing because of the END_OBSERVING_SETUP. Primary Mirror failed to make chase on first execution. Second time it ran. TRACK mode images of 5x2s, files luci#0172-0176.

after correcting END_OBSERVING_SETUP
07:39 execute Zibetti_PSB04_ImgJ.txt Files 0176 - 0179
07:49 execute Zibetti_PSB04_ImgH.txt Files 0180 - 0183
07:58 execute Zibetti_PSB04_ImgKs.txt Files 0184 - 0187

08:07 execute Zibetti_PSB04_STD_ImgKs.txt 08:10 execute Zibetti_PSB04_STD_ImgH.txt

Well...might as well get some experience...
08:16 execute Zibetti_PSB04_Acq.txt (spectroscopic acquisition).
08:17:01UT Error in sequence MOS - storage to turnout. WARNING hoding current still active at [card number 1, motor number 5...]
Called Volker, asked him to call Munich folks. Peter called, says it is fixed but my UI still shows red. UPDATE did not clear it. Restarted IMGUI but it is still red. Initialize cleared it. Lost 08:17 - 09:09 = 52m and now need to re-correct pointing & collimation. Put mask back (no mask in use).

09:18 Slew to pointing star ACT0381. We are having trouble with presets failing because M1 failed to make chase. Aaron reports that HP2 is having trouble (motor keeps turning but the mirror does not reach position). He is lowering and raising the mirror to watch the HP2 motor.

09:30 Thin cirrus coming in, getting thicker rapidly. No longer photometric.

09:30 Aaron called Chris. Aaron needs to go check the card on HP2 in the cell. Reseating the card did not help, and Aaron reports he can see the end of the motor shaft turning. Chris said this is indicating a mechanical issue and we cannot work more tonight. Lost remainder of the night: 3.5h. Issue #2448 was repaired during the following day.

10:30 We are done.

11:30 Taking darks. FLUSH_DETECTOR 3 times, then:
  • 10x2s int, NEXPO=5 (files 203-207)
  • 10x5s int, NEXPO=5 (files 208-212)
  • 10c10s int, NEXPO=5 (files 213-217)

Observing Plan

Closed dome
  • Test writing LUCIFER data to archive1 /newdata disk (/archive on LUCIFER computer)
  • Verify ingestion into archive, propagation to Tucson and on
  • Familiarization with LUCIFER UI

* Correct initial collimation/pointing with IE/CA

* Run dice5 pattern(s) to check that things still work as in May/June

* Check open-loop tracking performance (rms on off-axis GS vs. rms on LUCIFER) as reference data for comparison to post-Nov-HBS work and this has not been checked in a while.

* Collect astrometric field data to check telescope scale at colder temps

* Check guide star FWHM/rms vs LUCIFER image quality (closed loop DIQ)

* Make new collimation model, implement (optional if it looks good)
  • Derive and implement before pointing model obs

* Make new pointing model, implement (if coll model good)
  • Reset mount encoders (now takes only 2 minutes)
  • Have TO start the pointing log!
  • Check first position in AZ > 360 degrees wrap
  • Use ARC method to measure AGw/Rotator center
  • 28 points are sufficient, more is always better
  • Can use NB filter in LUCIFER, the pointing stars are ~7.5 mag
  • needs stable temps!
  • May need to do several short PMs as temps are changing rapidly

* Take Field aberration measurements if seeing is good (time-series, stability measurements)

* Demonstrate guiding to 18th mag (Tel=GUIDE)
* WFS on 18th mag. Will likely need to set the wfs max exposure time to 120s
* Quantify the guide/wfs limits with mag and seeing vs. LUCIFER DIQ

-- DavidThompson - 04 Dec 2009
Topic attachments
ISorted ascending Attachment Action Size Date Who Comment
daq.20091205.gifgif daq.20091205.gif manage 41 K 08 Dec 2009 - 23:39 JohnHill Plot of Telescope and Mirror Temperatures
pmcc.sx.thermal.20091205.gifgif pmcc.sx.thermal.20091205.gif manage 26 K 08 Dec 2009 - 23:39 JohnHill Detailed Plot of Primary Mirror Temperatures
psf.20091205.gifgif psf.20091205.gif manage 26 K 08 Dec 2009 - 23:40 JohnHill Plot of Primary Mirror Collimation
zern.20091205.gifgif zern.20091205.gif manage 17 K 08 Dec 2009 - 23:41 JohnHill Plot of Active Optics Corrections
Topic revision: r10 - 08 Dec 2009, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback